OpenStudy (anonymous):

Evaluate or simplify the numerical expressions. (16)^3/4

OpenStudy (anonymous):

this means the fourth root of 16 cubed take the fourth root of 16, cube the result

OpenStudy (anonymous):

How do you do that? That is what I'm stuck on. I don't know how to put it in my calculator. Can you plase explain it to me please?

OpenStudy (andriod09):

JEdwards85,\[16^{3/4}\] right? that's the problem?

OpenStudy (anonymous):

yes that is the problem

OpenStudy (andriod09):

okay, so first thing is first.16^ (-3/4) can be rewritten as: \[16^(3/4)= 1/(16^(1/4)^3)= 1/2^3 = 1/8= 0.125\] that's how its' solved. hope i helped!

OpenStudy (anonymous):

i did not see a minus sign in the exponent

OpenStudy (anonymous):

Thank you, that is okay

OpenStudy (anonymous):

the fourth root of 16 is 2, and you that because \(2^4=16\) not from a calculator \(2^3=8\) again by arithmetic

OpenStudy (andriod09):

it's a typo satilite73, nothing more.